# Unifying Bias and Unfairness in Information Retrieval: A Survey of Challenges and Opportunities with Large Language Models

### v1

@@ 6. Conclusion
This survey has delved into the emergence of new bias and unfairness challenges within IR systems in this LLM era. We have established a unified framework to understand these issues as distribution mismatch problems and systematically categorized mitigation strategies into data sampling and distribution reconstruction approaches. Through an in-depth review of several types of bias and unfairness, along with their corresponding mitigation strategies, we provide a comprehensive overview of the current progress. Furthermore, we have highlighted some open problems for future directions. We hope this survey can help researchers and stakeholders in the IR field and beyond to have a better understanding of bias and unfairness issues within IR systems in this LLM era, facilitating the development of more reliable and trustworthy IR systems.

### v2

@@ 5. Conclusion and Future Directions
This survey has delved into the emergence of new bias and unfairness challenges within IR systems in this LLM era. We have established a unified framework to understand these issues as distribution mismatch problems and systematically categorized mitigation strategies into data sampling and distribution reconstruction approaches. Through an in-depth review of fifteen types of bias and unfairness, along with their corresponding mitigation strategies, we provide a comprehensive overview of the current progress. Despite the considerable attention given to this topic, we identify some important problems for further exploration.
Biases and Unfairness in IR Feedback Loops. In real IR systems, the interaction between users, models, and information forms feedback loops that impact each other over time. These loops can significantly shape user perceptions and preferences based on the information they are exposed to. This interaction, in turn, influences the training data, creating a cycle that may reinforce existing biases and unfairness. Novel strategies to interrupt the feedback loops are essential for mitigating these issues.
Unified Mitigation Framework. Current methods primarily address individual instances of bias or unfairness, but in the future, we should consider unified solutions. This is because various types of bias and unfairness are not isolated. Presenting a unified framework can facilitate a deeper understanding of these relationships, enabling methods for addressing different types of bias and unfairness to complement each other. Our proposed unified perspective offers a potential direction to address these issues simultaneously.
Theoretical Analysis and Guarantees. The current exploration of bias and unfairness within the intersection between LLMs and IR systems has predominantly been through empirical studies. However, there is a critical need for robust theoretical analysis to augment these empirical findings. Future efforts should focus on developing more rigorous analytical frameworks.
Better Benchmarks and Evaluation. Most benchmarks currently utilized to study bias and unfairness within simulated environments. There is a crucial need for collecting large-scale, real-world datasets to enhance the evaluations and broaden research horizons. Additionally, as LLMs increasingly draw upon existing online data to train subsequent generations, dynamic benchmarks are needed to be constructed. Consequently, future work can focus on exploring a systematic evaluation protocol for different bias and unfairness issues.
Acknowledgements.
This work was funded by the National Key R&D Program of China (2023YFA1008704), the National Natural Science Foundation of China (No. 62377044, No.62276248), Beijing Key Laboratory of Big Data Management and Analysis Methods, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, PCC@RUC, funds for building world-class universities (disciplines) of Renmin University of China, and the Youth Innovation Promotion Association CAS under Grants No.2023111.